Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
hardwick, michimura, swinkels - 0:17 Thursday 21 June 2018 (41922) Print this report
Violin mode damping

Summary: Tonight we successfully implemented active violin mode damping. We demonstrated damping of modes on all test masses by length actuation on the marionettas. Modes that are isolated in frequency are easily damped into the noise floor. We had trouble with modes close together in frequency and on the same mirror, as expected; specifically around 449.3 Hz.

- - - - - - - -

Damping scheme

The damping scheme and associated channels are shown visually here. It goes as follows:

  1. take in DARM as error signal
  2. a BP filter is applied around violin mode: bandpass is 4th order buttorworth, ~200 mHz wide. There are two BP filters to chose from. Filter 1 is centered on measured violin frequency, Filter 2 is shifted by ~60 degrees
  3. apply gain to mode path
  4. all mode paths belonging to one mirror are summed together to create a mirror-specific correction signal
  5. correction signal is sent to either pitch or length of the marionetta

Damping notes

Gains and filters and drives that effectively damped are here. Times that damping started to occur are also there, in case anyone wants to look at data. All modes that are frequency-separated from other modes on same mirror were easily damped into the noise floor.

Ringdown activity here. Full before/after spectrum here. Bandpass coverage here. Note: we don't know what that broad bump in the spectrum is, but it's been there all night and we don't think its associated with our activity.

Only trouble modes are around 449.3 Hz. NE has modes at 449.30 and 449.33 Hz, making the sensing and independent damping difficult. We rang them up quite a bit before getting a setting that somewhat turned them around; they will need more work. Zoom in of spectrum around the modes here. Zoom of bandpass coverage mess here.

RMS channels are best indicators of damping/growing, but other nearby high modes can bleed in.

GUI

Yuta made an awesome GUI for violin mode damping, shown here!

For everything you can:

  • turn on/off a master switch. This is turned off by Metatron in the DOWN state, requires manual turn on (has 10 second ramp). In this way you can keep known good filters and gains set and eventually this can be turned on via Metatron.

For each mirror you can:

  • turn off the output
  • toggle between length/pitch marionetta drive
  • open relevant dataDisplay screens, which will start automatically.

For each mode you can:

  • get the current gain (wait time >5 sec)
  • set the gain
  • toggle between filter 1 & 2
  • emergency stop sending any signal along that mode path

GUI and associated config files are found in  /virgoDev/Automation/scripts/LSC/ViolinDamping/violingui.py. It can be opened for all mirrors ( >>python violingui.py ) and also for one mirror, input as an argument ( >>python violingui.py NE )

Other stuff

Violin damping infrustructure is in LSC_Acl.cfg, filters in LSC_Filters.cfg. Note: we hijacked ASC_{mirror}_TX/TY_NOISE_OUT channels to send to DSP for now. This will be cleaned up some, matrix options for drive, etc.

Average elapsed time of LSC_Acl increased form 12.8 to 16.6 us with all added filters, etc. We plan to next optimize some filters, etc., to bring this down.

Images attached to this report
Non-image files attached to this report
Comments to this report:
hardwick, michimura - 19:11 Friday 22 June 2018 (41949) Print this report

We've found filters and gains that damp all modes. All violin mode (1st harmonics) are damped to under 2e-21 in strain, shown in Hrec here (and zoomed here).

41949_20180622190636_violindampingdeforeafterwhole.png

Since gain values may need to be changed, we've keep the config default values as zero; the working values are stored in the GUI code. After a recompile, you can press "SET ALL gains" to reset to these values. The current working settings are also attached here. We note that only WI requires 60 deg phase correction (use filter 2). We haven't checked to see if all WI modes damp better with this.

At max, we are driving 0.02 V per correction signal.

Images attached to this comment
Non-image files attached to this comment
carbognani, michimura - 16:59 Tuesday 26 June 2018 (41997) Print this report
The violin mode damping GUI has been packaged as violinUI separate package and a production version (v1r1) has been uploaded in /virgoApp:
/virgoApp/violinUI/v1r1

This version has also been added under the IceWM menu SUS section and all declinations of the GUI can be started from there (Fig. 1)
Images attached to this comment
michimura - 14:07 Monday 06 August 2018 (42344) Print this report

I made a PyQt version of GUI for violing damping.
Previously, the GUI was made with Python Tkinter and current values are acquired using AclParamGet commants.
PyQt version uses ConnectOnline function to continuously get the current values (see logbook #42150 and comments).

PyQt version has almost the same functionality as before, but better since you can monitor the current values live and easier to understand the signal flow.
To save which filters are used and to save whether TX or TY is used, I added the following lines to LSC_SMS.cfg (this is the reason I had to restart LSC Acl; see logbook #42298).

ACL_SMS        "LSC_"        NI_VIOLIN1_ERR    NI_VIOLIN2_ERR    NI_VIOLIN3_ERR    NI_VIOLIN4_ERR    NI_VIOLIN5_ERR    NI_VIOLIN6_ERR    NI_VIOLIN7_ERR    NI_VIOLIN8_ERR
ACL_SMS        "LSC_"        NE_VIOLIN1_ERR    NE_VIOLIN2_ERR    NE_VIOLIN3_ERR    NE_VIOLIN4_ERR    NE_VIOLIN5_ERR    NE_VIOLIN6_ERR    NE_VIOLIN7_ERR    NE_VIOLIN8_ERR
ACL_SMS        "LSC_"        WI_VIOLIN1_ERR    WI_VIOLIN2_ERR    WI_VIOLIN3_ERR    WI_VIOLIN4_ERR    WI_VIOLIN5_ERR    WI_VIOLIN6_ERR    WI_VIOLIN7_ERR    WI_VIOLIN8_ERR
ACL_SMS        "LSC_"        WE_VIOLIN1_ERR    WE_VIOLIN2_ERR    WE_VIOLIN3_ERR    WE_VIOLIN4_ERR    WE_VIOLIN5_ERR    WE_VIOLIN6_ERR    WE_VIOLIN7_ERR    WE_VIOLIN8_ERR
ACL_SMS        "LSC_"        ASC_NI_TX_NOISE_CORR    ASC_NE_TX_NOISE_CORR    ASC_WI_TX_NOISE_CORR    ASC_WE_TX_NOISE_CORR
ACL_SMS        "LSC_"        ASC_NI_TY_NOISE_CORR    ASC_NE_TY_NOISE_CORR    ASC_WI_TY_NOISE_CORR    ASC_WE_TY_NOISE_CORR


"Restore Defaults" button on the top right will restore saved values of GAINs and filters and applies them.
Attached figure shows the current default GAINs and filters which were used in June 22 (at 13W; logbook #41949).

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×